OpenStudy (brittanydosey):

20. You saved $12,644.92 in an emergency fund. One fourth is in a regular savings account at a 3.5% APR, and the remainder is in a 30-day CD at a 4.57% APR. What is the difference in the amount of interest earned in 30 days between the two savings instruments? *

OpenStudy (brittanydosey):

Figure out how much is in each account. Then use the I = Prt formulas for the two different accounts. Remember, they give us time in days, not months. Once you have the two amounts, subtract them to get your final answer. @jackmullen55 @tylermcmullen23

OpenStudy (anonymous):

I=Prt is Percent* Rate* time

OpenStudy (anonymous):

i mean principle not percent

OpenStudy (anonymous):

I = 3161.23 × 0.035 × 0.08 = $ 8.85 + I = 9483.69 × 0.0457 × 0.08 = $ 34.67

OpenStudy (anonymous):

8.85 + 34.67

OpenStudy (brittanydosey):

43.52

OpenStudy (anonymous):

Yes thats right

OpenStudy (brittanydosey):

so that would be my answer? ?

OpenStudy (texaschic101):

12644.92 /4 = 3161.23 (savings account) 3.5% apr....= .035 I = 3161.23 * .035 * 30/365 I = 3161.23 * .035 * .08 I = 8.85 remainder 12644.92 - 3161.23 = 9483.69 (30 day CD) 4.57% apr = .0457 I = PRT I = 9483.69 * .0457 * .08 I = 34.67 difference : 34.67 - 8.85 = 25.82

OpenStudy (texaschic101):

difference means subtract